# From Equations to Insights: Unraveling Symbolic Structures in PDEs with LLMs

### v1

@@ 2.4 Inferencing
Then, the squared distance between two such operator sets, represented by vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as
$\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$
Clearly, it measures the number of mismatched operators between the two operator sets. Moreover, this metric prioritizes the presence of operators over their sequence order, ensuring that the predicted operators remain useful for numerical and analytical PDE solving, as well as for scientific analysis of PDE models.
3 Experimental Results
First, we analyze the ability of fine-tuned LLMs to predict the correct operator sets in PDE solutions. The objective is to assess whether the models can effectively learn and generalize the symbolic relationships between PDE solutions and problem data. We compare the predicted operator sets against ground-truth labels to measure the accuracy of the learned representations. Additionally, we examine the impact of model architecture by fine-tuning several commonly used open-source LLMs.
Second, we apply the fine-tuned LLMs to enhance the efficiency and accuracy of the FEX method by leveraging their predicted operator sets. This experiment attempts to achieve two goals. On one hand, we demonstrate that by obtaining more precise operator sets of smaller sizes, the computational costs spent in FEX can be significantly reduced. On the other hand, we evaluate whether FEX achieves improved accuracy in learning PDE solutions when provided with the correct prior information about the relevant operators. By comparing FEX solutions obtained with and without LLM-guided operator selection, we quantify improvements in both computational efficiency and numerical accuracy.
Through these experiments, we aim to establish the viability of fine-tuned LLMs as effective tools for discovering symbolic relations in PDE solutions and enhancing PDE-solving techniques.
3.1 Experimental Setup
For data generation, we consider two types of PDEs—namely, the Poisson equation and the linear conservation law—each paired with three commonly used boundary conditions: Cauchy, Dirichlet, and Neumann. Using a tree depth of 3, we randomly generate a dataset of 60,000 equations, with 30,000 examples for each PDE type, as described in Section 2.2. This dataset is then used to fine-tune two state-of-the-art large language models: BART (Lewis et al., 2019) and T5 (Raffel et al., 2020).
For fine-tuning, we optimize the cross-entropy loss using the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and a weight decay of $0.01$. The models are trained for 3 epochs with a batch size of 2. Fine-tuning is conducted on an NVIDIA A6000 GPU with 48 GB of memory. To efficiently manage storage, we save checkpoints every 500 steps and retain only the three most recent checkpoints.
Finally, we integrate the model into the FEX solver [^4] For a detailed description of the FEX method, we refer the reader to Liang and Yang (2022), and for the implementation of the original FEX, see https://github.com/LeungSamWai/Finite-expression-method.. During the execution of FEX for a given PDE, the model is invoked for inference. After inference, the post-processing step detailed in Section 2.4 is applied to extract the unique unary and binary operator sets. These sets are then used to configure the available operators within FEX during its search for an analytical PDE solution. We refer to this version of FEX as the "LLM-informed FEX" and denote the original FEX as "uninformed FEX." The uninformed FEX employs a fixed operator set:
[…]
3.2 Effectiveness of Fine-tuned LLMs
In this subsection, we evaluate the performance of the fine-tuned BART and T5 models in predicting operator sets for the two PDE types. Specifically, we track the training and validation losses throughout the training process, as well as the average number of mismatched operators on the test dataset for each epoch. The computational results are summarized in Tables 2 and 3.
Epoch Training Loss Validation Loss Average Mismatch
1 0.058200 0.057948 2.475833
2 0.040000 0.046193 2.020000
3 0.039200 0.038100 1.720210
Table 2: Fine-tuning BART on large dataset.
Epoch Training Loss Validation Loss Average Mismatch
1 0.046800 0.045303 1.711056
2 0.037200 0.036140 1.544056
3 0.012300 0.020392 1.012000
Table 3: Fine-tuning T5 on large dataset.
Both fine-tuned models exhibit a consistent decrease in training and validation loss across epochs, indicating successful fine-tuning. However, T5 demonstrates a more pronounced reduction in both loss metrics. The larger gap between BART’s training and validation loss suggests that it may be more prone to overfitting compared to T5, which appears to generalize better. In terms of the average number of mismatched operators, T5 consistently outperforms BART in this regard across all epochs. The sharper decline in mismatch for T5 suggests that it adapts more effectively to the symbolic representation of PDE solutions and captures the underlying operator structures with higher precision. Overall, T5 demonstrates superior performance in predicting operators for PDE solutions, making it a more suitable choice for integration into the LLM-informed FEX solver. In light of these findings, we incorporate the fine-tuned T5 model into the FEX framework to enhance its operator selection process.
3.3 LLM-informed FEX
In this subsection, we evaluate the effectiveness of the LLM-informed FEX using the fine-tuned T5 model in comparison to the uninformed FEX. As an initial validation step, we conduct a sanity check using several simple examples to ensure the correctness of the approach. To comprehensively assess the performance of the proposed framework, we further test both versions of FEX on a dataset of 100 randomly generated PDEs, consisting of 50 Poisson equations and 50 linear conservation law equations.
3.3.1 Sanity Check
We first construct 4 Poisson equations and 4 linear conservation law equations, whose true solutions, denoted as $u(x)$, are randomly handcrafted. For each example, we present the Binary and Unary set sizes, the number of iterations and computational time taken to find the approximate solution, and the error with respect to the true solution. The results are summarized in Tables 4 and 5, respectively.
$u(x)$ Method Binary Size Unary Size Iters Time [m] Error
$4\,\cos\bigl{(}4x^{2}\cos(x_{0})\bigr{)}$ LLM-informed 1 2 4.25 8.25 0
Uninformed 3 9 167 340 0
$4\,x_{1}^{3}+4\,x_{1}^{2}+2\,\cos\bigl{(}4\,x_{1}^{3}\cos(x_{0})\bigr{)}$ LLM-informed 2 4 102 286 $10^{-8}$
Uninformed [^5] Uninformed FEX did not converge within 2000 iterations, indicated by “2000+” (iterations) and “N/A” for error. 3 9 2000+ 2400+ N/A
$128\,x_{2}^{3}+2\,x_{2}\,e^{4\,x_{2}^{4}}$ LLM-informed 2 4 34.5 40.5 $4\times 10^{-7}$
Uninformed 3 9 90 186 $6\times 10^{-7}$
$64\,x_{1}^{2}\,e^{2\,x_{2}}$ LLM-informed 1 3 15.5 21 $3\times 10^{-8}$
Uninformed 3 9 103.5 161 $3\times 10^{-8}$
Table 4: Sanity check for the Poisson equation.
$u(x)$ Method Binary Size Unary Size Iters Time [m] Error
$\sin\Bigl{(}x_{1}+\tfrac{\pi}{4}x_{2}+\tfrac{\pi}{4}x_{3}\Bigr{)}$ LLM-informed 2 2 1 0.3 0
Uninformed 3 9 3 1 0
$4\,\bigl{(}2x_{0}-2x_{2}^{3}\bigr{)}^{3}$ LLM-informed 2 2 2 0.6 0
Uninformed 3 9 33.3 9.4 0
$8\,x_{2}^{2}+8\,\sin(x_{1})$ LLM-informed 2 3 1 1 0
Uninformed 3 9 24 28.5 0
[…]
Uninformed 3 9 273 488 0
Table 5: Sanity check for the linear conservation law euqation.
The computational results presented in Tables 4 and 5 demonstrate that the LLM-informed FEX significantly improves efficiency over the uninformed FEX while maintaining high accuracy. One major advantage is the reduction in the number of required operators. The uninformed FEX consistently uses a fixed operator set of 3 binary and 9 unary operators, whereas the LLM-informed FEX dynamically selects a smaller, more relevant subset—typically with at most 2 binary and 4 unary operators. This reduced search space leads to a dramatic decrease in the number of iterations required for convergence. For instance, in the first Poisson equation example, the LLM-informed FEX requires only 4.25 iterations compared to 167 for the uninformed version, and in another case, the uninformed FEX exceeds 2000 iterations without converging, while the LLM-informed approach successfully completes the solution in 102 iterations. This efficiency gain translates into significant computational time savings, with the LLM-informed approach solving problems up to an order of magnitude faster. Despite this drastic reduction in computational cost, the accuracy remains comparable; for cases where the uninformed FEX converges, both methods yield similar approximation errors, and for the linear conservation law equations, both achieve exact solutions. These results highlight the effectiveness of integrating the fine-tuned T5 model into the FEX framework, demonstrating that LLM-informed operator selection leads to a more efficient and reliable symbolic PDE solver.
3.3.2 Randomly Generated PDEs
To further validate the reliability of our approach, we evaluated the performance of the FEX algorithm on a set of randomly generated PDE examples.
For each PDE type, we randomly generated 50 instances using a depth-2 computational tree structure, as described in Section 2.2. Since FEX is a randomized reinforcement learning approach, each instance was solved five times to obtain average computational results, including the number of iterations until convergence, computation time, and solution error. The FEX algorithm was limited to a maximum of 150 iterations and an error tolerance of $10^{-3}$ (i.e., $\varepsilon<1\times 10^{-3}$), where $\varepsilon$ was computed during training using a composite loss that combines PDE residuals and boundary violations over a set of sampled points from the domain and its boundary, respectively. A PDE instance was considered successfully solved if the desired accuracy was achieved within 150 iterations. The computational results are summarized in Tables 6 and 7.
Table 6: Summary Statistics for Poisson PDEs
Metric Uninformed FEX LLM-Informed FEX Improvement (Reduction)
Success Ratio (Valid Solutions) 51.06% (24/47) 68.09% (32/47) $\uparrow 33.35\%$
For Cases Where BOTH Versions Yield Valid Solutions
Mean Iterations 61.80 27.87 $\downarrow 54.90\%$
Mean Duration (min) 40.73 20.25 $\downarrow 50.27\%$
Table 7: Summary Statistics for Linear Conservation Law PDEs
Metric Uninformed FEX LLM-Informed FEX Improvement (Reduction)
Success Ratio (Valid Solutions) 66.67% (30/45) 64.44% (29/45) $\downarrow 3.46\%$
For Cases Where BOTH Versions Yield Valid Solutions
Mean Iterations 24.87 13.98 $\downarrow 43.80\%$
Mean Duration (min) 6.43 3.60 $\downarrow 44.03\%$
The results in Tables 6 and 7 indicate that the LLM-informed FEX method consistently accelerates convergence compared to the Uninformed FEX, though its impact on solution validity varies across PDE types. For Poisson PDEs, the LLM-informed approach significantly improves the success ratio, increasing it by 33.35% (from 51.06% to 68.09%), while also reducing the number of iterations by 54.90% and computation time by 50.27%. This suggests that incorporating LLM-informed heuristics enhances both accuracy and efficiency for this class of problems. However, for linear conservation law PDEs, the LLM-informed approach slightly reduces the success ratio by 3.46%, indicating that its effectiveness in improving solution validity may be problem-dependent. This decline appears to stem from the inherent instability of FEX, even when accurate operators are provided. Despite this, the LLM-informed method still achieves a 43.80% reduction in iterations and a 44.03% decrease in computation time, demonstrating substantial gains in computational efficiency.
Overall, the LLM-informed approach consistently accelerates convergence, making it a valuable tool for solving PDEs more efficiently. However, its impact on solution validity remains inconsistent. This suggests that while the LLM-informed approach effectively reduce computational costs, one may need to conduct further refinement to the PDE solver itself to improve robustness across different PDE types.
4 Conclusions
In this work, we explored the potential of large language models (LLMs) to uncover symbolic relationships in partial differential equations (PDEs), a largely unexplored challenge in the intersection of AI and scientific computing. Our results demonstrate that fine-tuned LLMs can effectively predict the operators involved in PDE solutions by leveraging symbolic information from the governing equations. By integrating these predictions into the finite expression method (FEX), we significantly enhanced both the efficiency and accuracy of analytical PDE approximations. Compared to the traditional uninformed FEX, the LLM-informed approach reduces the number of operators required, accelerates convergence, and maintains high solution accuracy, providing a fully interpretable and computationally efficient pipeline. These findings highlight the promising role of LLMs in advancing symbolic reasoning for scientific problems, paving the way for further exploration of AI-driven methodologies in mathematical modeling and equation solving.
Acknowledgement

### v2

@@ C.1 Illustrative Example for Binary Encoding
The resulted mismatch can then be computed as $7$, indicating that the two functions are very different from each other.
C.2 Detailed Computational Results for Average Mismatch
Epoch LLaMA-8B LLaMA-3B T5 BART
1 0.82697 0.80510 1.71635 8.52542
2 0.65606 0.65040 1.40708 5.61399
3 0.55470 0.59328 1.37730 4.69332
4 0.51061 0.53747 1.37543 2.26791
5 0.48742 0.48601 1.32494 1.72084
6 0.45515 0.47934 1.36476 1.88165
7 0.42818 0.45894 1.34016 1.58241
8 0.41510 0.44808 1.32464 2.07886
9 0.39616 0.42293 1.31991 1.65313
10 0.39470 0.41894 1.30461 1.48738
11 0.38308 0.40672 1.30633 1.74097
12 0.38394 0.40697 1.29627 1.64541
Table 3: Comparison of Average Mismatch Across Models During Fine-tuning on 198K Dataset
C.3 Detailed computational results for FEX
True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $2\bigl{(}-2x_{2}^{3}+2\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl{(}4x_{2}\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl{(}2\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\exp(x_{1})+2\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\sin\bigl{(}2x_{3}^{3}-2\exp(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl{(}2x_{3}^{3}+2\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl{(}2x_{2}^{2}-2\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl{(}2\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl{(}2x_{1}^{2}-2x_{2}^{2}\bigr{)}$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl{(}2x_{3}^{4}-2\exp(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2x_{2}^{4}+2x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2x_{2}^{4}-2x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\sin\bigl{(}2\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl{(}2\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl{(}4x_{2}^{4}\bigr{)}\cos\bigl{(}2x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl{(}4\exp(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2x_{3}^{4}-2\bigr{)}^{3}+2\cos\bigl{(}2x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{4}+2\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\sin\bigl{(}2x_{1}^{4}-2\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\sin\bigl{(}2x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\cos\bigl{(}2\sin(x_{2})+2\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\cos\bigl{(}2x_{1}^{3}-2x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{2}+2\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$ $2\bigl{(}-2x_{2}^{3}+2\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl{(}4x_{2}\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl{(}2\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\exp(x_{1})+2\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\sin\bigl{(}2x_{3}^{3}-2\exp(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl{(}2x_{3}^{3}+2\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl{(}2x_{2}^{2}-2\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl{(}2\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl{(}2x_{1}^{2}-2x_{2}^{2}\bigr{)}$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl{(}2x_{3}^{4}-2\exp(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2x_{2}^{4}+2x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2x_{2}^{4}-2x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\sin\bigl{(}2\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl{(}2\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl{(}4x_{2}^{4}\bigr{)}\cos\bigl{(}2x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl{(}4\exp(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2x_{3}^{4}-2\bigr{)}^{3}+2\cos\bigl{(}2x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{4}+2\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$
Table 4: Results of LLM-informed and original FEX runs on 50 randomly generated Linear Conservation Law functions.
True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $8\,x_{1}\,\mathrm{exp}(x_{2})$ 42.4 1.2 74.44 0.5900 $3.75\times 10^{-4}$ $4.91\times 10^{-4}$ $2128\,x_{1}^{12}+$ 2 40.2 1.2 70.704 0.516 $3.97\times 10^{-2}$ $2.14\times 10^{-2}$ $-32\,x_{2}^{10}$ 35.2 1.2 71.586 0.536 $7.44\times 10^{-3}$ $6.71\times 10^{-3}$ $-8\,x_{1}^{4}\,x_{2}$ 25.0 1.2 53.342 0.510 $9.18\times 10^{-4}$ $3.86\times 10^{-4}$ $2\,\cos\bigl{(}4\,x_{2}^{6}\bigr{)}$ 44.2 3.0 85.792 4.326 $1.68\times 10^{-2}$ $1.23\times 10^{-2}$ $-8\,x_{2}\,\cos(x_{2})+2$ 32.4 2.4 78.076 4.548 $3.78\times 10^{-4}$ $3.06\times 10^{-4}$ $28\,x_{1}^{2}\,\cos(x_{2})+$ 2 42.8 3.2 85.55 6.154 $2.05\times 10^{-4}$ $1.85\times 10^{-4}$ $-4\,\mathrm{exp}(x_{3})+4\,\sin(x_{1})+2$ 20.2 1.6 43.712 2.150 $1.14\times 10^{-4}$ $2.04\times 10^{-4}$ 2($22\bigl{(}2\,x_{2}^{2}-2\,x_{3}^{4}\bigr{)}^{2}+$ 2 31.8 3.0 69.054 3.400 $6.55\times 10^{-3}$ $2.66\times 10^{-3}$ $-32\,\sin(x_{2})^{4}$ 15.6 1.6 36.210 1.704 $4.34\times 10^{-3}$ $4.66\times 10^{-3}$ $-4\,x_{2}^{3}-4\,x_{3}^{4}+16$ 20.6 2.4 45.110 2.716 $3.88\times 10^{-4}$ $4.68\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}-2\,\mathrm{exp}(x_{3})+2\,\cos(x_{3})\bigr{)}+$ 2 22.0 2.8 52.848 4.994 $3.83\times 10^{-4}$ $4.28\times 10^{-4}$ $28\,x_{1}^{2}\,x_{3}+$ 2 11.4 1.6 26.022 1.742 $3.58\times 10^{-4}$ $1.82\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{1})-2\,\cos(x_{3})\bigr{)}$ 19.2 3.0 47.168 4.768 $7.78\times 10^{-5}$ $3.13\times 10^{-5}$ $4\bigl{(}2\,\cos(x_{2})-2\,\cos(x_{3})\bigr{)}^{4}$ 10.2 1.6 21.852 1.476 $3.80\times 10^{-3}$ $3.12\times 10^{-3}$ 2($22\bigl{(}2\,x_{1}+2\,\mathrm{exp}(x_{1})\bigr{)}^{2}-$ 2 26.4 4.2 44.664 6.108 $1.07\times 10^{-3}$ $2.94\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{2}^{3}+2\,x_{3}^{4}\bigr{)}+$ 2 29.0 4.8 73.818 12.376 $6.12\times 10^{-4}$ $3.11\times 10^{-4}$ 2($22\bigl{(}-2\,x_{1}+2\,\mathrm{exp}(x_{3})\bigr{)}^{2}-$ 2 20.4 3.4 35.796 4.762 $1.96\times 10^{-3}$ $2.52\times 10^{-3}$ $22\,\cos\bigl{(}2\,\mathrm{exp}(x_{1})+2\,\sin(x_{2})\bigr{)}+$ 2 69.6 11.6 157.764 27.654 $3.92\times 10^{-5}$ $1.17\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{2})-2\,\cos(x_{2})\bigr{)}$ 18.2 3.2 37.792 5.254 $1.01\times 10^{-5}$ $3.35\times 10^{-5}$ $2\bigl{(}-2\,x_{2}^{3}+2\,\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\,\mathrm{exp}\bigl{(}4\,x_{2}\,\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8\,x_{2}^{4}-2\bigl{(}2\,\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4\,x_{2}^{4}+4\,x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\,\mathrm{exp}(x_{1})+2\,\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\,\sin\bigl{(}2\,x_{3}^{3}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4\,x_{2}^{4}+4\,x_{3}^{4}+8\,\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\,\sin\bigl{(}2\,x_{3}^{3}+2\,\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,x_{2}^{2}-2\,\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\,\sin(x_{2})^{2}\,\cos(x_{1})^{2}$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\,\mathrm{exp}\bigl{(}2\,\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\,\sin(x_{1})^{4}\,\cos(x_{1})^{4}$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\,\mathrm{exp}(2\,x_{2})-2\,\mathrm{exp}(2\,x_{1}^{2}-2\,x_{2}^{2})$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\,\mathrm{exp}\bigl{(}2\,x_{3}^{4}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2\,x_{2}^{4}+2\,x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2\,x_{2}^{4}-2\,x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2\,x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\,\sin\bigl{(}2\,\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\,\cos\bigl{(}4\,x_{2}^{4}\bigr{)}\,\cos\bigl{(}2\,x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\,\sin\bigl{(}4\,\mathrm{exp}(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2\,x_{3}^{4}-2\bigr{)}^{3}+2\,\cos\bigl{(}2\,x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\,\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{4}+2\,\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{1}^{4}-2\,\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\,\sin\bigl{(}2\,x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\,\cos\bigl{(}2\,\sin(x_{2})+2\,\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\,\cos\bigl{(}2\,x_{1}^{3}-2\,x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{2}+2\,\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232\,x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$
Table 5: Results of LLM-informed and original FEX runs on 50 randomly generated Poisson functions.